# Positive polytopes with few facets in the Grassmannian

### v1

@@ 4 Adjoints
The choice of degree in the definition above is explained as follows. Recall that the canonical bundle of $\mathbb{P}^{n}$ is $\mathcal{O}(-n-1)$ and that of $\mathrm{Gr}(2,4)$ is $\mathcal{O}(-4)$. Thus, any global rational differential top-form on $\mathbb{P}^{n}$ has degree $-n-1$, and any such form on $\mathrm{Gr}(2,4)$ has degree $-4$. For a polytope in $\mathbb{P}^{n}$ or a positive pentahedron/hexahedron in $\mathrm{Gr}(2,4)$ such a form is given by the ratio of the adjoint polynomial and the product of the facet equations. This means that for a polytope with $8$ facets in $\mathbb{P}^{5}$ the adjoint hypersurface is quadratic, and for a polytope with $7$ facets it is linear, i.e. a hyperplane. Analogously, a positive pentahedron in $\mathrm{Gr}(2,4)$ has a linear adjoint, and a positive hexahedron has a quadratic one.
@@ 4.1 Pentahedra
We now turn to characterizing $\mathcal{R}^{H}_{i}$. If $F$ is a face of $\mathrm{Gr}_{\geq 0}(2,4)$ that contributes a face to both $S_{1}$ and $S_{2}$, then it has to intersect $H$ inside $\mathrm{Gr}_{\geq 0}(2,4)$, and therefore $F\cap H$ does not contribute to either of the residual arrangements. If $F$ only contributes a face to $S_{1}$, then it is in $\mathcal{R}^{G}_{2}$ and the hyperplane spanned by $F$ intersects $H$ outside of $\mathrm{Gr}_{\geq 0}(k,n)$. This intersection is therefore in $\mathcal{R}^{H}_{1}$. If $F$ only contributes a face to $S_{2}$, then it already is in $\mathcal{R}^{G}_{1}$ and only contributes a lower-dimensional piece to $\mathcal{R}^{H}_{1}$. Thus, one sees that $\mathrm{span}(\mathcal{R}^{H}_{1})=\mathrm{span}(H\cap\mathcal{R}^{G}_{2})$ and, symmetrically, $\mathrm{span}(\mathcal{R}^{H}_{2})=\mathrm{span}(H\cap\mathcal{R}^{G}_{1})$. Since $\mathcal{R}_{i}=\mathcal{R}_{i}^{G}\cup\mathcal{R}_{i}^{H}$, we now can describe the span of $\mathcal{R}_{i}$, which is a projective space.
@@
where $b_{I}=0$ if $v_{I}\in\mathcal{R}_{1}$. The fact that the adjoint hyperplane should also interpolate the points $w_{j}$ imposes the conditions $b_{I}=c_{I}$ if $v_{I}\in S_{1}$ for the adjoint. This determines the equation of adjoint $a(\mathbf{p})$ uniquely.
@@ 4.2 Hexahedra
The variety $V$ consists of the $8$ positroid planes in $\mathrm{Gr}(2,4)$. In particular, each point of $V$ is contained in at least two positroid hyperplanes. Let $X$ be a component of $\mathcal{R}(\mathcal{P}_{S})$. Then, due to the observation above, $X\cap V$ is contained in the intersection of at least two facet hyperplanes of $S$. Moreover, we have $X\cap V\subseteq V\subseteq\mathrm{Gr}(2,4)$. Therefore, since $X\cap V$ is equal to the intersection of a subset of the facet hyperplanes of $\mathcal{P}_{S}$, $X\cap V$ is either in $S$ or in $\mathcal{R}(S)$. Since, due to convexity, $X$ is disjoint from $\mathcal{P}_{S}$ and $S=\mathcal{P}_{S}\cap\mathrm{Gr}(2,4)$, we conclude that $X\cap V$ is disjoint from $S$ and is therefore in $\mathcal{R}(S)$. Thus, $V\cap\mathcal{R}(\mathcal{P}_{S})\subseteq V\cap\mathcal{R}(S)$.
Conversely, let $Y$ be a component of $\mathcal{R}(S)$. Then $Y\cap V$ is contained in the intersection of facet hyperplanes of $S$. Each facet hyperplane of $S$ is also a facet hyperplane of $\mathcal{P}_{S}$, so $Y\cap V$ is either in $\mathcal{P}_{S}$ or in $\mathcal{R}(\mathcal{P}_{S})$. Since $S=\mathcal{P}_{S}\cap\mathrm{Gr}(2,4)$ and $Y\cap V\subseteq V\subseteq\mathrm{Gr}(2,4)$, $Y\cap V\not\subset S$, we conclude $Y\cap V\not\subset\mathcal{P}_{S}$ and thus $Y\cap V\subseteq\mathcal{R}(\mathcal{P}_{S})$. Thus, $V\cap\mathcal{R}(\mathcal{S})\subseteq V\cap\mathcal{R}(\mathcal{P}_{S})$.
@@
By Proposition 4.6, the number of quadrics in $I(\mathcal{R}(S))$ is bounded from above by the number of quadrics in $I(V\cap\mathcal{R}(\mathcal{P}_{S}))=\sqrt{I(V)+I(\mathcal{R}(\mathcal{P}_{S}))}$. We claim that as long as $h_{1}$ and $h_{2}$ are generic, the ideal $\sqrt{I(V)+I(\mathcal{R}(\mathcal{P}_{S}))}$ has at most seven linearly independent quadrics. The proof of this latter statement is a direct computation. The code is available at [14]. Since one of this quadrics is necessarily the defining polynomial of $\mathrm{Gr}(2,4)$, the family of adjoints is at most six-dimensional.
@@ 5 Examples
In this section we present examples of positive hexahedra in $\mathrm{Gr}(2,4)$ exhibiting different dimensions of the family of adjoints. In all of these examples there is however a unique canonical form with the required residues at the boundary vertices. We obtained these examples by using computer algebra systems. We identified the residual arrangements using the Mathematica implementation of cylindrical algebraic decomposition (see e.g. [4]), and the equations of adjoints were obtained in Macaulay2 [8]. We also verified that all positive hexahedra in our examples are connected and contractible, using the Julia package HypersurfaceRegions.jl presented in [5]. The details on our code are available at [14].
To analyze the possible canonical forms we further compute the dimension of adjoints to the algebraic boundary of one- and two-dimensional faces $F$ of the positive hexahedron. We write $\overline{F}$ for the Zariski closure of $F$. The intersection $\mathcal{R}(S)\cap\overline{F}$ may contain a divisor $D_{F}\subset\overline{F}$. The complement $(\mathcal{R}(S)\cap\overline{F})\setminus D_{F}$ is always contained in the residual arrangement $\mathcal{R}(F)$ of the face $F$. This containment may be strict. When two disjoint faces $F_{1}$ and $F_{2}$ have the same Zariski closure, i.e. $\overline{F_{1}}=\overline{F_{2}}$, then the residual arrangement of one face may contain boundary vertices of the other, while these vertices are not contained in the residual arrangement of $S$. We will show this phenomenon in Example 5.7. We note that there may be adjoints to $F$ on $\overline{F}$ that are not restrictions to $\overline{F}$ of adjoints to $S$. This will also be shown in Example 5.7.
We argue that each example we give of a positive hexahedron is a positive geometry, in the sense that it has a unique canonical form. In some cases, the positive hexahedron has a unique adjoint, while in other examples the adjoint is not unique, while the required ±$\pm 1$ residue at each boundary vertex is satisfied by a unique adjoint.
@@
where $\omega_{S}$ is a volume form, has iterated residues ±$\pm 1$ at the boundary vertices of every $1$-dimensional boundary face, then $(\mathrm{Gr}(2,4),S)$ is a positive geometry with canonical form $\Omega_{S}$.
@@
The form $\Omega_{S}$ has simple poles along the algebraic boundary of $S$, so its residue along each codimension one boundary component $F$ is a form $\Omega_{F}$ with iterated residues ±$\pm 1$ at its boundary vertices. Since the hyperplane arrangement is simple, the form $\Omega_{F}$ has simple poles along its boundary. Now, if $\Omega_{F}$ is not unique, there are two forms whose iterated residues at the boundary vertices coincide. There difference is then a form $\Omega^{\prime}_{F}$ with simple poles along the boundary of $F$ that vanishes in all boundary points. But then the iterated residues of $\Omega^{\prime}_{F}$ along $1$-dimensional faces also vanish. Inductively, the residue of $\Omega^{\prime}_{F}$ along its simple codimension $1$ boundary components must vanish while it has simple poles along this boundary, so $\Omega^{\prime}_{F}$ itself vanishes. We then get that the iterated residue of $\Omega_{S}$ along each face is the unique form with simple poles along the codimension one boundary of the face and iterated residues ±$\pm 1$ at its boundary vertices. Thus, the lemma follows.
As in [10, Lemma 2.16] we find a unique canonical form on a finite collection of disjoint intervals on the real line.
@@
Consider an $n$-dimensional space of rational $1$-forms on $\mathbb{P}^{1}$, each with $22n-$ 2 zeros and with simple poles in $n$ ordered pairs of real points $(p_{i},q_{i}),i=1..n$ so that $p_{1}<q_{1}<\ldots<p_{n}<q_{n}$ with respect to the chosen orientation. If the forms in this space have no common zeros, then there is a unique form with residue $1$ at each $p_{i}$ and $-1$ at each $q_{i}$. In particular, this form is the canonical form of the disjoint union of intervals $(p_{i},q_{i}),i=1,\ldots,n$.
@@
where for each $s$ in a $n$-dimensional vector space, $f_{s}$ is a polynomial of degree $2(n-2)$. The form $\Omega_{s}$ may be decomposed as a sum of $n$ forms by the partial fraction decomposition:
@@
This form has residues $c_{s,i}$ at $p_{i}$ and $-c_{s,i}$ at $q_{i}$, for $i=1,...,n$. The forms $\Omega_{s}$ have no common zeros, so for a general $s$ the scalars $c_{s,i}$ are all nonzero. The vector space of forms $\Omega_{s}$, is $n$-dimensional so there is a unique $s=s_{0}$, such that $c_{s_{0},i}=1$ for all $i$.
@@
The residual arrangement $\mathcal{R}(S)$ of $S$ is now given by six lines and a quadric surface:
$\{p_{12}p_{34}-p_{13}p_{24}+p_{14}p_{23}=h_{1}(\mathbf{p})=h_{2}(\mathbf{p})=0\},$
$\{p_{12}=p_{13}=p_{23}=h_{1}(\mathbf{p})=0\},\quad\{p_{13}=p_{23}=p_{34}=h_{1}(\mathbf{p})=0\},$
$\{p_{12}=p_{14}=p_{24}=h_{2}(\mathbf{p})=0\},\quad\{p_{14}=p_{24}=p_{34}=h_{2}(\mathbf{p})=0\},$
$\{p_{12}=p_{13}=p_{23}=p_{34}=0\},\quad\{p_{12}=p_{14}=p_{24}=p_{34}=0\}.$
There is a unique adjoint given by the polynomial $a_{S}=2p_{12}^{2}+p_{12}p_{13}+16p_{12}p_{14}+5p_{13}p_{14}+5p_{12}p_{23}+4p_{12}p_{24}+32p_{13}p_{24}+7p_{23}p_{24}-34p_{12}p_{34}+2p_{13}p_{34}+8p_{14}p_{34}+10p_{23}p_{34}+2p_{24}p_{34}+2p_{34}^{2}$.
@@
Let us also see how the positive hexahedron $S$ in this example is a $4$-dimensional positive geometry.
We check the iterated residues at the boundary vertices along each $1$-dimensional face, and conclude by Lemma 5.1. In what follows, we write $e_{ij}$ for the point in $\mathbb{P}^{5}$ whose vector of homogeneous coordinates satisfies $p_{ij}=1$ and $p_{kl}=0$ for $(k,l)\neq(i,j)$.
@@
The residual arrangement of the positive hexagon $S=\mathrm{Gr}(2,4)\cap\mathcal{P}$ is given by eight lines, a conic and a point. The lines and the conic are defined by the equations
$\{p_{23}=h_{1}(\mathbf{p})=h_{2}(\mathbf{p})=p_{12}p_{34}-p_{13}p_{24}=0\},$
$\{p_{12}=p_{13}=p_{23}=h_{1}(\mathbf{p})=0\},\quad\{p_{13}=p_{23}=p_{34}=h_{1}(\mathbf{p})=0\},$
$\{p_{12}=p_{13}=p_{14}=h_{1}(\mathbf{p})=0\},\quad\{p_{13}=p_{14}=p_{34}=h_{1}(\mathbf{p})=0\},$
$\{p_{23}=p_{24}=p_{34}=h_{2}(\mathbf{p})=0\},\quad\{p_{12}=p_{23}=p_{24}=h_{2}(\mathbf{p})=0\},$
$\{p_{12}=p_{14}=p_{24}=p_{34}=0\},\quad\{p_{12}=p_{13}=p_{14}=p_{34}=0\}.$
The two points in $\{qq=p_{12}=p_{34}=h_{1}=h_{2}=0\}$ have nonnegative coordinates, except $p_{14}$ and $p_{24}$ that are connected by the equation $011p_{14}^{2}-46p_{14}p_{24}-13p_{24}^{2}=$ 0. Projectively this equation has two solutions, in one of which $p_{14}$ and $p_{24}$ have different signs. So one of these points belongs to $S$, while the other one is a point in the residual arrangement.
@@
Next, there are three points on the quadric surface $\{qq=h_{1}=h_{2}=0\}$ of intersection between the conic plane sections $\{p_{12}p_{14}p_{34}=0\}$. There are four points in the quadric surface $\{qq=p_{12}=p_{34}=0\}$, the coordinate point $e_{14}$ and three points of intersection between the conic plane sections $\{p_{23}h_{1}h_{2}=0\}$. Furthermore, there are six points on the quadric surface $\{qq=p_{14}=p_{23}=0\}$, the coordinate points $e_{12},e_{34}$ and two points on each of the conic plane sections $\{h_{1}h_{2}=0\}$, so that the four plane sections $\{p_{12}h_{1}p_{34}h_{2}=0\}$ meet cyclically in boundary vertices. Finally two boundary vertices are the endpoints of a boundary segment on the line $\{p_{13}=p_{14}=p_{24}=h_{2}\}$ defined by $\{p_{12}p_{24}=0\}$.
There are $7$ boundary segments on conics, and $21$ boundary segments on lines, with one segment on each curve. Each boundary vertex is a boundary point of four segments. The one-skeleton of $S$ is connected, so as in the previous example, the iterated residue of the form
@@
along each boundary segment has no zeros and two simple poles. By the connectedness of the one-skeleton one may scale $a_{S}$ so that the iterated residues of $\Omega_{S}$ at the boundary vertices are ±$\pm 1$. By Lemma 5.1, $(\mathrm{Gr}(2,4),S)$ is a positive geometry with canonical from $\Omega_{S}.$
@@
$\{p_{14}=p_{23}=qq=0\},$
$\{p_{12}=p_{13}=p_{23}=p_{24}=0\},$
$\{p_{13}=p_{23}=p_{24}=p_{34}=0\},$
$\{p_{12}=p_{13}=p_{14}=p_{24}=0\},$
$\{p_{13}=p_{14}=p_{24}=p_{34}=0\},$
$\{p_{12}=p_{34}=h_{1}(\mathbf{p})=h_{2}(\mathbf{p})=0\}.$
The space of quadrics interpolating the residual arrangement $\mathcal{R}(S)$ of $S$ is spanned by

### v2

@@ 4 Adjoints
The choice of degree in the definition above is explained as follows. Recall that the canonical bundle of $\mathbb{P}^{n}$ is $\mathcal{O}(-n-1)$ and that of $\mathrm{Gr}(2,4)$ is $\mathcal{O}(-4)$. Thus, any global rational differential top-form on $\mathbb{P}^{n}$ has degree n1-n-1, and any such form on $\mathrm{Gr}(2,4)$ has degree 4-4. For a polytope in $\mathbb{P}^{n}$ or a positive pentahedron/hexahedron in $\mathrm{Gr}(2,4)$ such a form is given by the ratio of the adjoint polynomial and the product of the facet equations. This means that for a polytope with $8$ facets in $\mathbb{P}^{5}$ the adjoint hypersurface is quadratic, and for a polytope with $7$ facets it is linear, i.e. a hyperplane. Analogously, a positive pentahedron in $\mathrm{Gr}(2,4)$ has a linear adjoint, and a positive hexahedron has a quadratic one.
@@ 4.1 Pentahedra
We now turn to characterizing $\mathcal{R}^{H}_{i}$. If $F$ is a face of $\mathrm{Gr}_{\geq 0}(2,4)$ that contributes a face to both $S_{1}$ and $S_{2}$, then it has to intersect $H$ inside $\mathrm{Gr}_{\geq 0}(2,4)$, and therefore $F\cap H$ does not contribute to either of the residual arrangements. If $F$ only contributes a face to $S_{1}$, then it is in $\mathcal{R}^{G}_{2}$ and the hyperplane spanned by $F$ intersects $H$ outside of $\mathrm{Gr}_{\geq 0}(2,4)$. This intersection is therefore in $\mathcal{R}^{H}_{1}$. If $F$ only contributes a face to $S_{2}$, then it already is in $\mathcal{R}^{G}_{1}$ and only contributes a lower-dimensional piece to $\mathcal{R}^{H}_{1}$. Since each face $F$ of $\mathrm{Gr}_{\geq 0}(2,4)$ contributes a face to etiher $S_{1}$ or $S_{2}$, and since all the irreducible components of $\mathcal{R}_{i}^{H}$ are by definition Zariski closures of the sets of the form $F\cap H$ for some face $F$ of $\mathrm{Gr}_{\geq 0}(2,4)$, one sees that $\mathrm{span}(\mathcal{R}^{H}_{1})=\mathrm{span}(H\cap\mathcal{R}^{G}_{2})$ and, symmetrically, $\mathrm{span}(\mathcal{R}^{H}_{2})=\mathrm{span}(H\cap\mathcal{R}^{G}_{1})$. Since $\mathcal{R}_{i}=\mathcal{R}_{i}^{G}\cup\mathcal{R}_{i}^{H}$, we now can describe the span of $\mathcal{R}_{i}$, which is a projective space.
@@
where $b_{I}=0$ if $v_{I}\in\mathcal{R}_{1}$. The fact that the adjoint hyperplane should also interpolate the points $w_{j}$ imposes the conditions $b_{I}=Ac_{I}$ if $v_{I}\in S_{1}$ for the adjoint, where $A$ is some scalar factor. This is because each point $w_{j}$ is of the form $\alpha v_{I}+\beta v_{J}$ for the vertices $v_{I}$ and $v_{J}$ in $\mathcal{R}_{G}^{2}$. In addition, for any hyperplane interpolating $w_{j}$ we have $0=\alpha b_{I}+\beta b_{J}$, and, since $H$ contains $w_{j}$, we also have $0=\alpha c_{I}+\beta c_{J}$ for all $I$ and $J$ such that $v_{I},v_{J}\in\mathcal{R}_{G}^{2}$. This determines the equation of the adjoint $a(\mathbf{p})$ uniquely up to a scalar factor. This scalar factor is then fixed by computing the residue of the resulting canonical form at any vertex of $S_{1}$ and requiring it to be $\pm 1$ depending on the chosen orientation.
@@ 4.2 Hexahedra
The variety $V$ consists of the $8$ positroid planes in $\mathrm{Gr}(2,4)$, that is, of the linear two-dimensional components of the algebraic boundary of $\mathrm{Gr}_{\geq 0}(2,4)$. In particular, each point of $V$ is contained in at least two of the four positroid hyperplanes $\{p_{12}=0\}$, $\{p_{14}=0\}$, $\{p_{23}=0\}$, $\{p_{34}=0\}$. Let $X$ be a component of $\mathcal{R}(\mathcal{P}_{S})$. Then, due to the observation above and because $X\cap V\subseteq V\subseteq\mathrm{Gr}(2,4)$, each component of $X\cap V$ is contained in the intersection of Zariski closures of at at least two facets of $S$. Thus, since each component of $X\cap V$ is equal to the intersection of a subset of the facet hyperplanes of $\mathcal{P}_{S}$, $X\cap V$ is either in $S$ or in $\mathcal{R}(S)$. Since, due to convexity, $X$ is disjoint from $\mathcal{P}_{S}$ and $S=\mathcal{P}_{S}\cap\mathrm{Gr}(2,4)$, we conclude that $X\cap V$ is disjoint from $S$ and is therefore in $\mathcal{R}(S)$. Thus, $V\cap\mathcal{R}(\mathcal{P}_{S})\subseteq V\cap\mathcal{R}(S)$.
Conversely, let $Y$ be a component of $\mathcal{R}(S)$. Then $Y\cap V$ is contained in the intersection of facet hyperplanes of $S$. Each facet hyperplane of $S$ is also a facet hyperplane of $\mathcal{P}_{S}$, so $Y\cap V$ is either in $\mathcal{P}_{S}$ or in $\mathcal{R}(\mathcal{P}_{S})$. Since $S=\mathcal{P}_{S}\cap\mathrm{Gr}(2,4)$ and $Y\cap V\subseteq V\subseteq\mathrm{Gr}(2,4)$, and $Y\cap V\not\subset S$, we conclude $Y\cap V\not\subset\mathcal{P}_{S}$ and thus $Y\cap V\subseteq\mathcal{R}(\mathcal{P}_{S})$. Thus, $V\cap\mathcal{R}(\mathcal{S})\subseteq V\cap\mathcal{R}(\mathcal{P}_{S})$.
@@
By Proposition 4.6, the number of quadrics in $I(\mathcal{R}(S))$ is bounded from above by the number of quadrics in $I(V\cap\mathcal{R}(\mathcal{P}_{S}))=\sqrt{I(V)+I(\mathcal{R}(\mathcal{P}_{S}))}$. We claim that as long as $h_{1}$ and $h_{2}$ are generic, the ideal $\sqrt{I(V)+I(\mathcal{R}(\mathcal{P}_{S}))}$ has at most seven linearly independent quadrics. The proof of this latter statement is a direct computation. The code is available at [14]. Since one of these quadrics is necessarily the defining polynomial of $\mathrm{Gr}(2,4)$, the family of adjoints is at most six-dimensional.
@@ 5 Examples
In this section we present examples of positive hexahedra $S$ in $\mathrm{Gr}(2,4)$ exhibiting different dimensions of the family of adjoints. In all of these examples there is however a unique canonical form with the required residues at the boundary vertices. We obtained these examples by using computer algebra systems. We identified the residual arrangements using the Mathematica implementation of cylindrical algebraic decomposition (see e.g. [4]), and the equations of adjoints were obtained in Macaulay2 [8]. We also verified that all positive hexahedra in our examples are connected and contractible, using the Julia package HypersurfaceRegions.jl presented in [5]. The details on our code are available at [14].
To analyze the possible canonical forms we consider one- and two-dimensional faces $F$ of the positive hexahedron $S$. They are semialgebraic sets with a boundary consisting of lower dimensional faces. We write $\overline{F}$ for the Zariski closure of $F$ and consider the arrangement ${\mathcal{H}}_{F}$ in $\overline{F}$ of Zariski closures of codimension one boundary faces in $F$. The residual arrangement $\mathcal{R}(F)$ is the collection of intersections in $\overline{F}$ of codimension at least two of Zariski closures of these boundary faces. Let ${\mathcal{O}}(K_{F})$ be the canonical bundle on $\overline{F}$. Similar to adjoints of $S$ in $\mathrm{Gr}(2,4)$, an adjoint hypersurface or adjoint divisor of $F$ contains the residual arrangement $\mathcal{R}(F)$ and is the zero locus of a global section of ${\mathcal{O}}(K_{F}+{\mathcal{H}}_{F})$.
The intersection $\mathcal{R}(S)\cap\overline{F}$ may contain a divisor $D_{F}\subset\overline{F}$. The complement $(\mathcal{R}(S)\cap\overline{F})\setminus D_{F}$ is always contained in the residual arrangement $\mathcal{R}(F)$ of the face $F$. This containment may be strict. When two disjoint faces $F_{1}$ and $F_{2}$ have the same Zariski closure, i.e. $\overline{F_{1}}=\overline{F_{2}}$, then the residual arrangement of one face may contain boundary vertices of the other, while these vertices are not contained in the residual arrangement of $S$. We will show this phenomenon in Example 5.8. We note that there may be adjoints to $F$ on $\overline{F}$ that are not restrictions to $\overline{F}$ of adjoints to $S$. This will also be shown in Example 5.8.
We argue that each example we give of a positive hexahedron is a positive geometry, in the sense that it has a unique canonical form. In some cases, the positive hexahedron has a unique adjoint, while in other examples the adjoint is not unique, while the required $\pm 1$ residue at each boundary vertex is satisfied by a unique adjoint.
@@
where $\omega_{S}$ is a volume form, has iterated residues $\pm 1$ at the boundary vertices of every $1$-dimensional boundary face, then $(\mathrm{Gr}(2,4),S)$ is a positive geometry with canonical form $\Omega_{S}$.
@@
The form $\Omega_{S}$ has simple poles along the algebraic boundary of $S$, so its residue along each codimension one boundary component $F$ is a form $\Omega_{F}$ with iterated residues $\pm 1$ at its boundary vertices. Since the hyperplane arrangement is simple, the form $\Omega_{F}$ has simple poles along its boundary. The adjoint $a_{S}$ is unique, so the lemma follows if the form $\Omega_{F}$ is unique for each $F$.
Now, if $\Omega_{F}$ is not unique, there are two forms whose poles and iterated residues at the boundary vertices coincide. Their difference is then a form $\Omega^{\prime}_{F}$ with simple poles along the boundary of $F$ that vanishes in all boundary points. But then the iterated residues of $\Omega^{\prime}_{F}$ along one-dimensional faces also vanish. Inductively, the residue of $\Omega^{\prime}_{F}$ along its simple codimension $1$ boundary components must vanish while it has simple poles along this boundary, so $\Omega^{\prime}_{F}$ itself vanishes and $\Omega_{F}$ is unique.
We then get that the iterated residue of $\Omega_{S}$ along each face is the unique form with simple poles along the codimension one boundary of the face and iterated residues $\pm 1$ at its boundary vertices. Thus, the lemma follows.
As in [10, Lemma 2.16] there is a unique canonical form on a finite collection of disjoint intervals on the real line. The following lemma provides a criterion for a space of rational one-forms to contain this canonical form.
@@
Consider an $n$-dimensional space of rational one-forms on $\mathbb{P}^{1}$, each with 2n22n-2 zeros and with simple poles in $n$ ordered pairs of real points $(p_{i},q_{i}),i=1..n$ so that $p_{1}<q_{1}<\ldots<p_{n}<q_{n}$ with respect to the chosen orientation. If the forms in this space have no common zeros, then there is a unique form with residue $1$ at each $p_{i}$ and 1-1 at each $q_{i}$. In particular, this form is the canonical form of the disjoint union of intervals $(p_{i},q_{i}),i=1,\ldots,n$.
@@
where for each $s$ in an $n$-dimensional vector space, $f_{s}$ is a polynomial of degree 2n22n-2. The form $\Omega_{s}$ may be decomposed as a sum of $n$ forms by the partial fraction decomposition:
@@
This form has residues $c_{s,i}$ at $p_{i}$ and $c_{s,i}$ at $q_{i}$, for i=1,,ni=1,...,n. The forms $\Omega_{s}$ have no common zeros, so for a general $s$ the scalars $c_{s,i}$ are all nonzero. The vector space of forms $\Omega_{s}$, is $n$-dimensional so there is a unique $s=s_{0}$, such that $c_{s_{0},i}=1$ for all $i$.
Remark 5.5.
If the space of one-forms in the lemma has a common zero, it may not contain the canonical form.
@@
The residual arrangement $\mathcal{R}(S)$ of the corresponding positive hexahedron $S$ in $\mathrm{Gr}(2,4)$ is now given by six lines and a quadric surface:
$Q:=\{p_{12}p_{34}-p_{13}p_{24}+p_{14}p_{23}=h_{1}(\mathbf{p})=h_{2}(\mathbf{p})=0\},$
$L_{1}:=\{p_{12}=p_{13}=p_{23}=h_{1}(\mathbf{p})=0\},\quad L_{2}:=\{p_{13}=p_{23}=p_{34}=h_{1}(\mathbf{p})=0\},$
$L_{3}:=\{p_{12}=p_{14}=p_{24}=h_{2}(\mathbf{p})=0\},\quad L_{4}:=\{p_{14}=p_{24}=p_{34}=h_{2}(\mathbf{p})=0\},$
$L_{5}:=\{p_{12}=p_{13}=p_{23}=p_{34}=0\},\quad L_{6}:=\{p_{12}=p_{14}=p_{24}=p_{34}=0\}.$
[…]
Figure 2: Intersections between the components of the residual arrangement in Example 5.6
The incidence relations between the components of this residual arrangement are depicted in Figure 2. There is a unique adjoint given by the polynomial $a_{S}=2p_{12}^{2}+p_{12}p_{13}+16p_{12}p_{14}+5p_{13}p_{14}+5p_{12}p_{23}+4p_{12}p_{24}+32p_{13}p_{24}+7p_{23}p_{24}-34p_{12}p_{34}+2p_{13}p_{34}+8p_{14}p_{34}+10p_{23}p_{34}+2p_{24}p_{34}+2p_{34}^{2}$.
@@
Let us also see how the positive hexahedron $S$ in this example is a four-dimensional positive geometry.
We check the iterated residues at the boundary vertices along each one-dimensional face, and conclude by Lemma 5.1. In what follows, we write $e_{ij}$ for the point in $\mathbb{P}^{5}$ whose vector of homogeneous coordinates satisfies $p_{ij}=1$ and $p_{kl}=0$ for $(k,l)\neq(i,j)$.
@@
As before $\mathrm{Gr}(2,4)$ is defined in Plücker coordinates by $\{qq:=p_{12}p_{34}-p_{13}p_{24}+p_{14}p_{23}=0\}$. The residual arrangement of the positive hexagon $S=\mathrm{Gr}(2,4)\cap\mathcal{P}$ is given by eight lines, a conic and a point. The lines and the conic are defined by the equations
$C_{1}:=\{p_{23}=h_{1}(\mathbf{p})=h_{2}(\mathbf{p})=p_{12}p_{34}-p_{13}p_{24}=0\},$
$L_{1}:=\{p_{12}=p_{13}=p_{23}=h_{1}(\mathbf{p})=0\},\quad L_{2}:=\{p_{13}=p_{23}=p_{34}=h_{1}(\mathbf{p})=0\},$
$L_{3}:=\{p_{12}=p_{13}=p_{14}=h_{1}(\mathbf{p})=0\},\quad L_{4}:=\{p_{13}=p_{14}=p_{34}=h_{1}(\mathbf{p})=0\},$
$L_{5}:=\{p_{23}=p_{24}=p_{34}=h_{2}(\mathbf{p})=0\},\quad L_{6}:=\{p_{12}=p_{23}=p_{24}=h_{2}(\mathbf{p})=0\},$
$L_{7}:=\{p_{12}=p_{14}=p_{24}=p_{34}=0\},\quad L_{8}:=\{p_{12}=p_{13}=p_{14}=p_{34}=0\}.$
The two points in $\{qq=p_{12}=p_{34}=h_{1}=h_{2}=0\}$, have nonnegative coordinates, except $p_{14}$ and $p_{24}$ that are connected by the equation $11p_{14}^{2}-46p_{14}p_{24}-13p_{24}^{2}=0$. Projectively this equation has two solutions, in one of which $p_{14}$ and $p_{24}$ have different signs. So one of these points belongs to $S$, while the other one is a point in the residual arrangement. The incidence relations between the one-dimensional components of this residual arrangement are depicted in Figure 3. The residual point is disjoint from all one-dimensional components.
[…]
Figure 3: Intersections between the components of the residual arrangement in Example 5.7
@@
Next, we find boundary vertices on two-dimensional faces of $S$ in quadric surfaces. There are three points on the quadric surface $\{qq=h_{1}=h_{2}=0\}$ of intersection between the conic plane sections $\{p_{12}p_{14}p_{34}=0\}$. There are four points on the quadric surface $\{qq=p_{12}=p_{34}=0\}$, the coordinate point $e_{14}$ and three points of intersection between the conic plane sections $\{p_{23}h_{1}h_{2}=0\}$. Furthermore, there are six points on the quadric surface $\{qq=p_{14}=p_{23}=0\}$, the coordinate points $e_{12},e_{34}$ and two points on each of the conic plane sections $\{h_{1}h_{2}=0\}$, so that the four plane sections $\{p_{12}h_{1}p_{34}h_{2}=0\}$ meet cyclically in boundary vertices. Finally two boundary vertices are the endpoints of a boundary segment on the line $\{p_{13}=p_{14}=p_{24}=h_{2}\}$ defined by $\{p_{12}p_{34}=0\}$.
Altogether there are seven boundary segments on conics, and $21$ boundary segments on lines, with one segment on each curve. Each boundary vertex is a boundary point of four segments. The one-skeleton of $S$ is connected, so as in the previous example, the iterated residue of the form
@@
along each boundary segment has no zeros and two simple poles. By the connectedness of the one-skeleton one may scale $a_{S}$ so that the iterated residues of $\Omega_{S}$ at the boundary vertices are $\pm 1$. By Lemma 5.1, $(\mathrm{Gr}(2,4),S)$ is a positive geometry with canonical from $\Omega_{S}.$
@@
$Q:=\{p_{14}=p_{23}=qq=0\},$
$L_{1}:=\{p_{12}=p_{13}=p_{23}=p_{24}=0\},$
$L_{2}:=\{p_{13}=p_{23}=p_{24}=p_{34}=0\},$
$L_{3}:=\{p_{12}=p_{13}=p_{14}=p_{24}=0\},$
$L_{4}:=\{p_{13}=p_{14}=p_{24}=p_{34}=0\},$
$\{p_{12}=p_{34}=h_{1}(\mathbf{p})=h_{2}(\mathbf{p})=qq=0\}.$
@@
Figure 4: Intersections between the components of the residual arrangement in Example 5.8
The two points are disjoint from the higher-dimensional components, whose incidence relations are in turn depicted in Figure 4. The space of quadrics interpolating the residual arrangement $\mathcal{R}(S)$ of $S$ is spanned by